Conside the following linear programming problem

Maximize $4X1 + $8X2
Subject To 2X1 + 5X2 ≤50

Constraint A

3X1 + 3X2 ≤ 48

Constraint B


X1, X2 ≥ 0

Constraint C
Answer #1

Corner points Z=4x1+8x2
0,0 4(0)+8(0)=0+0=0
0,10 4(0)+8(10)=80
10,6 4(10)+6(10)=40+60=100
0,16 4(0)+16(10)=160